OpenStudy (anonymous):

You invest an initial $2,000 in an account that has an annual interest rate of 6%, compounded daily. How much money will you have in the account after 15 years? Round your answer to the nearest whole number.

OpenStudy (anonymous):

Apply this formula: \[A = P \left( 1 + \frac{ r }{ n } \right) ^{nt}\] * A = final amount * P = principal amount (initial investment) * r = annual nominal interest rate (as a decimal, not in percentage) * n = number of times the interest is compounded per year * t = number of years